OpenStudy (anonymous):

Suppose you were given f(x) = x²+x-6 and g (x) = x-2. What if you had to calculate f(x) times g(x). What would you do? What is the product?

OpenStudy (phi):

I think they want you to multiply the two polynomials together

OpenStudy (anonymous):

First use the distributive property: \[ (x^2+x-6) \times (x-2) = x(x^2+x-6)-2(x^2+x-6) \]Do it again \[ x^3+x^2-6x-2x^2-2x+12 \]Add like terms (same exponents) \[ x^3-x^2-8x+12 \]